OpenStudy (anonymous):

Julie is a math tutor in the academic center. She charges $210 for the first 7 hours of tutoring and $20 for each additional hour of tutoring. If she earned $410 tutoring 1 student, how many total hours did she spent tutoring the student?

OpenStudy (anonymous):

A. h = 7 × 410 + 210 20 B. h = 7 − 410 + 210 20 C. h = 7 − 410 − 210 20 D. h = 7 + 410 − 210 20

OpenStudy (anonymous):

I could help you, but it would be $20 an hour...

OpenStudy (anonymous):

wow

OpenStudy (netlopes1):

410=210+20x -> 20x=200 -> x=10, or 10h.

OpenStudy (anonymous):

17

OpenStudy (netlopes1):

Sorry, i am wrong, i don't see "7" h.

OpenStudy (netlopes1):

So, make 10h (first resolution) + 7h (additions) -> 17h

OpenStudy (anonymous):

To get a feel for these types of problems, sometimes it's best to just try different hours and see how much she got. Obviously, there's an algebraic way to do it (as netlopes has explained) but the algebraic way is meaningless if you don't understand what's going on. So... In 7 hours, she earns 210 dollars. In 8 hours, she earns 210 plus 20 more, so 230. In 9 hours, she earns 210 plus 40 more, so 260. Notice that, after the first 7 hours, we're just going up by 20 every hour. So how many times can I go up by 20 to get from 210 to 410? Well, that's 200 dollars, and 200 divided by 20 is 10...so it must be 10 extra hours. Add that to the seven original hours and you get 17. Now, you just have to figure out which of your answers is equal to 17.